# : A Benchmark for Data-Efficiency in Long-Horizon Indoor Mobile Manipulation Robotics

### v5

@@ III-B Simulated Dataset
The dataset for our simulated tasks consists of 521 trajectories collected across 20 rooms in five distinct environments, designed to ensure diversity in both object types and room layouts. We utilized the AI2THOR [20] simulator and specifically employed RoboTHOR [21] environments, which feature multiple interconnected rooms suitable for long-horizon tasks.
For robot control, we used ManipulaTHOR [22], which includes an arm capable of fine manipulation and provides low-level pose data. Each trajectory includes a range of attributes, including but not limited to RGB-D egocentric observations, segmentations, the poses of the robot body, end-effector, and grasped objects, as well as discrete actions taken during the tasks.
To generate realistic and diverse instructions, we leveraged Prolific111prolific.com for crowdsourcing to gather 521 natural language commands from 41 participants, reflecting tasks that people would naturally expect robots to perform in indoor settings. Using these commands, we teleoperated the robot in the simulator with a custom keyboard-based teleoperation system we developed, ensuring accuracy and consistency throughout the data collection process. After teleoperation, each trajectory was replay-verified to confirm correctness and eliminate errors, ensuring high-quality data integrity.
@@ III-C Real-World Dataset
The dataset for our real-world tasks comprises 50 trajectories collected across 10 rooms in three distinct environments. These environments were chosen for their large size, structural features that allow for the creation of long-horizon tasks, and their status as workplaces where robots can be particularly useful for daily activities. The first environment is a three-room laboratory, the second is a floor in a university building, and the third spans two adjacent floors connected by stairs within the same building. These spaces include kitchens, furnished lobbies, a classroom, and a staircase, offering diverse objects and room layouts to enhance the generalizability of trained models.
To handle complex terrain and ensure stability, we selected the Spot [23] quadruped mobile manipulator, equipped with an arm for pick-and-place tasks. The quadruped’s design enables it to navigate stairs and maintain balance during task execution, making it ideal for the challenging settings of our chosen environments.
@@
We teleoperated the Spot robot using a tablet equipped with thumbsticks to collect task trajectories based on 50 natural language commands provided by seven participants. The commands, distributed as 20, 15, and 15 across the three environments, specified long-horizon room-to-room and floor-to-floor pick-and-place tasks that reflect practical scenarios in workplace settings. Task trajectories were collected at 3 Hz to ensure detailed and realistic data.
IV Evaluation
IV-A Baseline Models
To evaluate the and establish baselines for future benchmarking, we test both learning-based and neuro-symbolic models. Given the growing popularity of imitation learning (IL), particularly BC, we benchmark two state-of-the-art (SOTA) BC models. These models provide insight into how data-efficient current BC algorithms are by being tested on , validating that the dataset supports meaningful learning by demonstrating that models can achieve measurable yet limited success. Additionally, we benchmark a neuro-symbolic system that integrates foundation models with task and motion planning (TAMP). This system, evaluated zero-shot without requiring training, investigates whether non-BC methods can achieve greater data efficiency while achieving comparable or superior performance, offering guidance for future development of data-efficient MoMa methodologies. Finally, we assess the performance of a random agent, wherein actions are uniformly sampled to establish a lower-bound baseline.
RT-1. Robotics Transformer (RT-1) [4] is a BC model designed for MoMa, leveraging a Transformer architecture to process natural language instructions and RGB image inputs into discretized actions. Its training on MoMa tasks, along with its status as an open-source, SOTA learning-based MoMa model, makes it particularly suitable for benchmarking on as a baseline.
To adapt RT-1 for the ManipulaTHOR robot, we modified its action space to align with the robot’s capabilities. For task navigation, the robot base operates using discrete actions, while continuous deltas (discretized into 256 bins) are utilized for body rotation and manipulation to achieve precise control.
MotionGlot-MoMa MG [24] is designed to address the problem of varying action and observation dimensions in multi-embodiment models [25]. Leveraging training principles from language models, MG proposes a training procedure for motion generation across multiple embodiments as a next-token prediction problem. We adapt MotionGlot (MG) for MoMa tasks by leveraging ’s demonstrations consisting of paired RGB observations and actions, calling it MG-MoMa. MG-MoMa is trained to learn a language-conditioned model that maps those observations to corresponding actions. This process involves two stages: First, a VQ-VAE is employed to learn a discrete latent codebook for the images, forming the observation vocabulary, while discrete binning is applied to tokenize the actions. In the second stage, both the observation and action vocabularies are integrated with the text vocabulary from GPT-2 [26]. This unified vocabulary allows for action generation similar to text generation. Finally, the instruction-tuning template from MG-MoMa is used for text command-based action generation.
LIMP. Language Instruction grounding for Motion Planning (LIMP) [27] is a neuro-symbolic system for pick-and-place MoMa that integrates Linear Temporal Logic (LTL), foundation models, and novel TAMP algorithms to process natural language instructions into executable TAMP plans. By translating natural language instructions into LTL formulas enriched with Composable Referent Descriptors, LIMP dynamically generates context-aware specification translations and compiles them into finite-state automata for verifiably correct execution. Unlike learning-based methods, LIMP requires no training, allowing it to be evaluated zero-shot. This enables a comparison of its data efficiency against BC models, providing insight into whether such systems are a promising direction for developing more data-efficient approaches to MoMa.
@@ IV-A Baseline Models
Data Split & Generalization. To evaluate generalization capabilities, we conduct experiments using carefully designed data splits for scene generalization and task generalization on the simulation dataset, highlighted in Table II.
@@
For scene generalization, we employ a 5-fold cross-validation approach, where each of the five scenes in the dataset is held out as the test scene in turn. The held-out test set consists of all trajectories from the designated scene, while the training and validation sets are derived from the remaining four scenes. Specifically, 80% of the trajectories from each of the four remaining scenes are included in the training set, and 20% are allocated to the validation set, following a stratified split to preserve data distribution. During each fold, we train and hyperparameter tune a separate model specific to the held-out test scene. This process ensures that we assess the model’s ability to generalize to entirely unseen environments while maintaining experimental controls. The rotation of held-out scenes allows for a comprehensive evaluation of scene generalization across the entire dataset.
@@
For task generalization, the data split is consistent across all five scenes. The training set comprises 80% of the trajectories from each scene, the validation set includes 10%, and the test set includes the remaining 10%. Unlike scene generalization, the held-out test set in this experiment evaluates the model’s ability to generalize to novel tasks rather than unseen environments. Hyperparameter tuning is performed, resulting in a single model optimized for task generalization on the test set.
@@
Training. To benchmark the data efficiency of modern BC methods, we train the RT-1 and MotionGlot models from scratch without any pretrained parameters on the simulation tasks. This approach evaluates how well these models and their architectures learn from our modestly sized yet realistic dataset, providing a measure of their data efficiency specifically for long-horizon MoMa tasks. Both models are trained using the scene generalization and task generalization experiments, ensuring their capabilities are evaluated under varying conditions.
@@
Fine-tuning. To further evaluate data efficiency, we fine-tune RT-1 by loading pretrained parameters and testing its performance on the simulation tasks. Empirical analysis showed that freezing some model layers led to better performance than not freezing any layers, so this approach was used for the final experiment. Using this setup, we conducted the task generalization experiment to measure how much pretrained MoMa knowledge impacts RT-1’s ability to adapt efficiently to new MoMa tasks.
V Discussion
V-A Results
Scene Generalization. The success rates for the Scene Generalization experiment, detailed in Table III, are exceptionally low, with an overall average of just and across all scenes for RT-1 and MG-MoMa respectively. None of the individual scenes achieved success rates exceeding , indicating significant challenges in generalizing to novel scenes with unseen room layouts, furniture configurations, and object placements for long-horizon MoMa tasks. Both models outperform the random agent, demonstrating their ability to learn from the data and satisfy the minimum performance criterion.
Scenes 1 and 3 exhibited relatively higher success rates compared to the others for RT-1. This can be attributed to a relatively high proportion of evaluation tasks in these scenes involving objects that were positioned directly in front of or adjacent to the robot’s initial spawn location, reducing the complexity of navigation and manipulation required. In contrast, Scene 1 for RT-1 exhibited the lowest performance, likely due to the robot spawning in an enclosed and isolated room where many tasks necessitated navigating out of the room to retrieve the target object, significantly increasing task difficulty.
Interestingly, for MG-MoMa, Scenes 1 and 3 were the exact opposite of RT-1, with those being the lowest performers. It is unclear why that is the case, but it could be that MG-MoMa fails to recognize nearby objects. Similarly, Mg-MoMa’s best performers are RT-1’s worst, Scenes 2 and 5. Overall, MG-MoMa demonstrates performance that is near RT-1, which is noteworthy given that Motionglot was not originally intended for mobile manipulation tasks.
Task Generalization. The success rates, presented in Table IV, for the Task Generalization experiment for RT-1 and Mg-MoMa, are low overall, though they are slightly higher compared to the average performance in the Scene Generalization experiment. This improvement is likely attributed to the model having access to all the scenes during training, enabling it to better generalize to new tasks within already familiar environments. Additionally, while it is unsurprising that RT-1 outperformed MG-MoMa on unseen demonstrations, it is notable that the two models performed comparably on seen demonstrations, and the margin of RT-1’s outperformance was relatively small. This is remarkable, particularly given that MotionGlot was originally designed for cross-embodied motion generation rather than specifically for mobile manipulation.
We also evaluated on seen tasks, which were uniformly sampled from the training data to match the size of the unseen data from the test split to measure performance on previously encountered tasks. Surprisingly, RT-1 achieved a marginally higher success rate on unseen tasks () compared to seen tasks (). This indicates that the model’s training did not overfit to the training data, which is a positive outcome. One plausible explanation is that the sampler may have inadvertently selected a disproportionate number of especially challenging tasks for the seen evaluation set, rendering them overly difficult for the model despite having been part of the training data.
Task Generalization with Fine-Tuning. Interestingly, the unseen task performance was marginally lower compared to the fully trained model’s performance on unseen tasks, as seen in Table IV. This result suggests the pretrained weights may have skewed the model’s representation toward the original training tasks and away from effectively adapting to the unique challenges of our benchmark, especially since the original was trained in the real world.
LIMP. The results presented thus far indicate that learning-based models, face significant challenges in addressing the complexities of long-horizon MoMa tasks. To explore an alternative paradigm, we benchmarked the LIMP system on the Task Generalization experiment.
The evaluation was conducted in real-world environments using Spot, assessing 45 tasks, including a mix of tasks originally specified for both simulated and real-world environments. Since LIMP was originally designed for real-world execution, this mixed evaluation approach was necessary to also test its performance on simulation-specified tasks. For these simulation-designed tasks, real-world setups were constructed to approximate their simulated counterparts. This enabled us to evaluate LIMP’s versatility across both task types, providing insights into its adaptability to different environmental contexts. As shown in Table IV, LIMP achieved a relatively impressive average success rate of on unseen tasks, vastly outperforming RT-1 and MG-MoMa without any training, making it relatively more data efficient.
One key factor contributing to LIMP’s success is its use of pre-generated maps of the environments. They provide the robot with precise information about the absolute and relative locations of objects and regions, eliminating the need to search and enabling it to interpret the fine-grained semantics specified in language-conditioned tasks. Additionally, the maps allow the robot to utilize robust motion planning algorithms for navigation, which are less susceptible to covariate shift compared to the mapless autoregressive action decoding used in BC models like RT-1. This robustness is critical for the long-horizon navigation aspect of the tasks, where the inherent large observation and state spaces exacerbate the challenges faced by learning-based methods.
Another significant advantage of LIMP is its incorporation of multimodal foundation models, which endow it with extensive prior knowledge to reason about task requirements and perceptual inputs in a zero-shot manner. This capability enables LIMP to handle the semantic and contextual complexities of our benchmark tasks without requiring task-specific training. While LIMP demonstrates considerable promise, it still falls short of perfect task execution, indicating room for improvement in handling the intricacies of our long-horizon MoMa tasks. Overall, LIMP represents a promising direction for future system development, blending the strengths of learning-based approaches with those of classical planning.
V-B Ablations
Impact of Dataset Size. The poor performance observed with learning-based methods is expected to stem from the limited, realistically sized dataset of the benchmark, but this needs to be further examined. Will simply scaling the data improve the performance? Will these models effectively and efficiently use more data to learn better? To address this, we investigate how performance scales when progressively larger subsets of the dataset (, , , and ) are used for training. This evaluates whether these models can effectively utilize limited data or require substantially more, directly testing their data efficiency and scalability for long-horizon MoMa tasks.
The ablation results from the Task Generalization setup, summarized in Table V, surprisingly reveal no clear upward trend in performance gains despite progressively increasing the training data for RT-1 and hyperparameter tuning per subset. We attribute part of the variation to the inherent complexity of these long-horizon MoMa tasks and the variability in the sampled subsets, where smaller subsets can incidentally include a higher proportion of comparatively easier scenarios. Overall, this suggests that simply providing more samples does not guarantee better learning outcomes for these tasks, and it is likely that limited data is not the primary cause of the poor performance. Moreover, since the linear increase in data size did not result in proportional improvement, it indicates that the model is unable to efficiently learn from the additional data provided, deeming it data inefficient. This raises questions about deeper limitations, such as architectural constraints, warranting further investigation.
@@ V-B Ablations
| MG-MoMa | – | – | ||
| LIMP | – | – | – | |
| Random | – | – | ||
Impact of Architecture. To determine if data inefficiency stems from the architecture used, we conduct a model comparison experiment by evaluating a different architecture. While MotionGlot was already benchmarked and provides some architectural comparison with RT-1, it is not sufficiently different since it also relies on a Transformer architecture. To explore a more distinct alternative, we developed a custom LSTM-based model by replacing RT-1’s Transformer with an LSTM while retaining the same encoders for a consistent and isolated comparison, as LSTMs have been previously used for sequential robotic tasks. The LSTM variant is evaluated under the task generalization setup using subset experiments with , , , and of the dataset, enabling us to analyze how its performance scales with increasing data. By comparing the full dataset results and the performance rate of change (RoC) against RT-1, we aim to determine whether data inefficiency is tied to architectural limitations or if the issue lies elsewhere.
As shown in Table V, the LSTM variant exhibits no upward trend when moving from to of the data, indicating it is also data inefficient. To compare the trends between the two architectures, we analyze their average rates of change. Both exhibit similarly non-positive trends, indicating that the data inefficiency is unlikely to primarily stem from the choice of architecture and seems to span across different architectures, since they are not efficiently making use of more data. Furthermore, since RT-1’s performance is somewhat better than the LSTM variant’s, and its RoC is slightly higher, it indicates marginally better data efficiency, further supporting the argument that architecture is unlikely the primary cause of RT-1’s data inefficiency observed in the main experiments.
Since architecture does not appear to be the primary issue, it is plausible that these long-horizon indoor MoMa tasks are simply inherently too complex and challenging for learning-based methods. Identifying the exact causes of data inefficiency in these models requires further investigation, which is beyond the scope of this work and is left for the community to explore in future work.

### v6

@@ III-B Simulated Dataset
The dataset for our simulated tasks consists of trajectories, with an average length of timesteps, collected across rooms in five distinct environments, designed to ensure diversity in both object types and room layouts. We utilized the AI2THORÂ [28] simulator, specifically the RoboTHORÂ [29] environments, which feature connected rooms suitable for long-horizon tasks.
For robot control, we used ManipulaTHORÂ [30], which includes an arm capable of fine manipulation and provides low-level pose data. Each trajectory includes a range of attributes, including but not limited to RGB-D egocentric observations, segmentations, the poses of the robot body, end-effector, and grasped objects, as well as discrete actions.
To generate realistic and diverse instructions, we leveraged Prolific222prolific.com for crowdsourcing to gather natural language commands from participants, reflecting tasks that people would naturally expect robots to perform in indoor settings. We teleoperated the simulated robot with a custom keyboard-based teleoperation system we developed to execute these commands. After teleoperation, each trajectory was replay-verified to confirm correctness, ensuring data integrity.
@@ III-C Real-World Dataset
The dataset for our real-world tasks comprises trajectories collected across rooms in three distinct environments. These environments were chosen for their large size, structural features that allow for the creation of long-horizon tasks, and their status as workplaces where robots can be particularly useful for daily activities. The first environment is a three-room laboratory, the second is a floor in a university building, and the third spans two adjacent floors connected by stairs within the same building. These include kitchens, furnished lobbies, a classroom, and a staircase, offering diverse objects and room layouts to challenge model generalizability.
@@
To handle these complex terrains, we teleoperated the Spot robot using a tablet equipped with thumbsticks to collect task trajectories based on natural language commands provided by seven participants. The commands, distributed as , , and across the three environments, respectively, specified long-horizon room-to-room and floor-to-floor pick-and-place tasks that reflect practical scenarios in workplace settings. The trajectories were collected at Hz.
@@
IV Evaluation
IV-A Baseline Models
To evaluate and establish baselines for future benchmarking, we test both learning-based and neuro-symbolic systems. Given the growing popularity of imitation learning, particularly BC, we benchmark two state-of-the-art (SOTA) BC models. This provides insight into the data efficiency of current BC algorithms.
Additionally, we benchmark a neuro-symbolic system that integrates foundation models with task and motion planning (TAMP). The foundation model modules are trained on large-scale visual and textual data, but the system overall requires no robot demonstration training. Thus, we evaluated it zero-shot to assess whether non end-to-end methods can achieve greater data efficiency. By comparing its performance to learning-based approaches, this evaluation provides insight into whether such methods can achieve comparable or superior results, offering guidance for future development of data-efficient MoMa methodologies. Finally, we assess the performance of a random agent to establish a lower-bound baseline, wherein actions are uniformly sampled.
RT-1. Robotics Transformer (RT-1) [4] is a BC model designed for MoMa, leveraging a transformer architecture [31] to process natural language instructions and RGB image inputs into discretized actions. Its training on MoMa tasks, along with its status as an open-source, SOTA learning-based MoMa model, makes it particularly suitable for benchmarking on as a baseline.
To adapt RT-1 for the ManipulaTHOR robot, we modified its action space to align with the robotâs capabilities. For task navigation, the robot base operates using discrete actions, while continuous deltas (discretized into 256 bins) are utilized for body rotation and manipulation.
MotionGlot-MoMa. MotionGlot (MG) [32] addresses the problem of varying action and observation dimensions in multi-embodiment models [33]. Leveraging training principles from language models, MG proposes a training procedure for motion generation across multiple embodiments as a next-token prediction problem.
We adapt MG for MoMa by leveraging âs demonstrations consisting of paired RGB observations and actions, calling it MG-MoMa. MG-MoMa is trained to learn a language-conditioned model that maps those observations to corresponding actions. This process involves two stages: First, a VQ-VAEÂ [34] is employed to learn a discrete latent codebook for the images, forming the observation vocabulary, while discrete binning is applied to tokenize the actions. In the second stage, both the observation and action vocabularies are integrated with the text vocabulary from GPT-2Â [35]. This unified vocabulary enables action generation akin to text generation, utilizing the instruction-tuning template from MG-MoMa for text command-based action execution.
LIMP. Language Instruction grounding for Motion Planning (LIMP) [36] is a neuro-symbolic system for pick-and-place MoMa integrating Linear Temporal Logic (LTL), Large Language Model (LLM), VLM, and novel TAMP algorithms to process natural language instructions into executable TAMP plans. By translating instructions into LTL formulas enriched with Composable Referent Descriptors [36], LIMP dynamically generates context-aware specification translations and compiles them into finite-state automata for verifiably correct execution. Unlike learning-based methods, LIMP requires no training on robot demonstrations, allowing it to be evaluated zero-shot. This enables a comparison of its data efficiency against BC models, providing insight into whether such systems are a promising direction for developing more data-efficient approaches to MoMa.
@@ IV-A Baseline Models
To evaluate data efficiency, we assess how well models generalize to novel scenarios using Scene Generalization and Task Generalization, two widely used experimental setups [4, 15]. These evaluations are essential for benchmarking data efficiency because they test whether models can learn effectively from limited data and apply that knowledge to new scenarios. Scene Generalization reflects real-world deployment challenges where robots must perform tasks in previously unseen environments with different layouts and/or objects, while Task Generalization assesses adaptability to new tasks within seen settings. Together, these experiments quantify a modelâs data efficiency by isolating generalization challenges at both the scene and task levels, providing insights into how well models leverage limited training data to achieve robust performance.
@@
Data Split & Generalization. To implement these evaluations, we conduct experiments using carefully designed data splits on the simulation dataset, highlighted in Table II. For Scene Generalization, we employ a -fold cross-validation approach, where each of the five scenes in the dataset is held out as the test scene in turn. The held-out test set consists of all trajectories from the designated scene, while the training and validation sets are derived from the remaining four scenes. Specifically, of the trajectories from each of the four remaining scenes are included in the training set, and are allocated to the validation set, following a stratified split to preserve data distribution. During each fold, we train and hyperparameter tune a separate model specific to the held-out test scene. This process ensures that we assess the modelâs ability to generalize to entirely unseen environments while maintaining experimental controls. The rotation of held-out scenes allows for a comprehensive evaluation of Scene Generalization across the entire dataset.
@@
For Task Generalization, the data split is consistent across all five scenes. The training set comprises of the trajectories from each scene, the validation set includes , and the test set includes the remaining . Unlike Scene Generalization, the held-out test set in this experiment evaluates the modelâs ability to generalize to novel tasks rather than unseen environments. Hyperparameter tuning is performed, resulting in a single model optimized for Task Generalization on the test set.
@@
Training. To benchmark the data efficiency of modern BC methods, we train the RT-1 and MG-MoMa models from scratch without any pretrained parameters on the simulation tasks. This approach evaluates how well these models learn from our modest yet realistically sized dataset, providing a measure of their data efficiency specifically for long-horizon MoMa tasks. Both models are trained using the Scene Generalization and Task Generalization experiments, ensuring their capabilities are evaluated under varying conditions.
@@
Fine-Tuning. We fine-tune RT-1 and MG-MoMa with pretrained parameters and evaluate their performance on the simulation tasks. Empirical analysis showed that for RT-1, fine-tuning with some frozen model layers led to better performance than not freezing any, so this approach was used for the final experiment. In contrast, for MG-MoMa, the unfrozen model performed better, leading to the use of fully trainable parameters. Using this setup, we conducted the Task Generalization experiment to measure how much pretrained MoMa knowledge impacts the modelsâ ability to adapt efficiently to new MoMa tasks.
Metric. To evaluate model performance, we use a task success rate metric that quantifies the completion of long-horizon MoMa tasks. Each task consists of four sequential subtasks: (1) navigating to the target object, (2) grasping the object, (3) transporting the object to the goal location, and (4) placing the object at the goal. Each successful subtask contributes one point, with a maximum of four points per task. The success rate is calculated by averaging the subtask scores within each task and then averaging across all tasks. This metric captures both navigation and manipulation performance, providing insight into a modelâs ability to execute multi-step long-horizon tasks.
@@
V Discussion
V-A Results
Scene Generalization. The success rates for the Scene Generalization experiment, detailed in Table III, are exceptionally low, with an overall average of just and across all scenes for RT-1 and MG-MoMa respectively. This demonstrates significant challenges in generalizing to novel scenes with unseen room layouts, furniture configurations, and object placements for our long-horizon MoMa tasks. Both models outperform the random agent, demonstrating at least a minimal capacity to learn something from the data.
Scenes 1 and 3 exhibited relatively higher success rates compared to the others for RT-1. This can be attributed to a relatively high proportion of evaluation tasks in these scenes involving objects that were positioned directly in front of or adjacent to the robotâs initial spawn location, reducing the complexity of navigation and manipulation required. In contrast, Scene 2 for RT-1 exhibited the lowest performance, likely due to the robot spawning in an enclosed and isolated room where many tasks necessitated navigating out of the room to retrieve the target object, significantly increasing task difficulty. Overall, MG-MoMa demonstrates performance that is near RT-1, which is noteworthy given that MG was not originally intended for MoMa tasks.
Task Generalization. Table IV presents the results of our Task Generalization experiments, which assess the modelâs ability to adapt to new tasks in a familiar environments. Notably, both RT-1 and MG-MoMa show slightly higher overall performance compared to the Scene Generalization experiments. This improvement is likely due to the larger training dataset size (see Table II) and the availability of all scenes during training. Moreover, while the performance of the baselines is equal for the seen settings, RT-1 outperforms MG-MoMa by on the unseen environments. Meanwhile, the random agent showed no difference in performance between seen and unseen tasks, as expected.
Task Generalization with Fine-Tuning. Interestingly, RT-1âs unseen task performance was marginally lower than the fully trained modelâs, as shown in Table IV. Similarly, MG-MoMa showed lower performance on seen tasks. This may mean that the pretrained parameters biased the models toward their original training tasks and away from effectively adapting to the unique challenges of , especially since the original models were trained in the real world.
@@ V-A Results
| MG-MoMa | ||||
| LIMP | â | â | â | |
| Random | â | â | ||
LIMP. The results presented thus far show that learning-based models (BC in particular) face significant challenges in addressing the complexities of âs long-horizon MoMa tasks. To explore an alternative paradigm, we benchmarked the LIMP system on the Task Generalization experiment.
The evaluation was conducted in real-world environments using Spot, assessing tasks, including a mix of tasks originally specified for both simulated and real-world environments. Since LIMP was originally designed for real-world execution, this mixed evaluation approach was necessary to also test its performance on simulation-specified tasks. For these simulation-designed tasks, real-world setups were constructed to approximate their simulated counterparts. This enabled us to evaluate LIMPâs versatility across both task types, providing insights into its adaptability to different environmental contexts. As shown in Table IV, LIMP achieved a relatively impressive average success rate of on unseen tasks, vastly outperforming RT-1 and MG-MoMa without any training on robot demonstrations, making it relatively more data efficient.
One key factor contributing to LIMPâs success is its ability to generate instruction-conditioned semantic maps of the environment. These maps provide the robot with precise information about the locations of objects and regions of interest, eliminating the need to search and enabling the interpretation of fine-grained semantics specified in language-conditioned tasks. Additionally, the maps allow the robot to utilize robust motion planning algorithms for navigation, which are less susceptible to covariate shift compared to the mapless action decoding used in BC models like RT-1. This robustness is critical for the tasksâ long-horizon navigation aspect, where the inherent large observation and state spaces exacerbate the challenges faced by learning-based methods.
Another significant advantage of LIMP is its incorporation of multimodal foundation models, which endow it with extensive world knowledge to reason about task requirements and perceptual inputs in a zero-shot manner. This capability enables it to handle the semantic and contextual complexities of our benchmark tasks without requiring task-specific training. While it demonstrates considerable promise, it still falls short of perfect execution, indicating room for improvement in handling the intricacies of our long-horizon MoMa tasks. Overall, LIMP represents a promising direction for future MoMa development, blending the strengths of learning-based approaches with those of classical planning.
V-B Ablations
Impact of Dataset Size. ALOHA Unleashed [37], a bimanual manipulation system trained on K demonstrations, only achieved a success rate of for certain tasks, showing that even with large datasets, high performance is not guaranteed. Similarly, Lin et al. [38] found that increasing demonstrations generally improves manipulation performance, though with diminishing returns. Since our benchmark involves MoMaâwith larger action and observation spaces due to added navigationâand uses far fewer trajectories, substantial performance improvements likely require more data. To investigate this, we pose the following questions: Are the baseline models data inefficient? Will scaling the data improve their performance? To answer these, we investigate how performance scales when progressively larger subsets of the simulation dataset (, , , and ) are used for training in the Task Generalization experiment. This evaluates whether these models can effectively utilize limited data or require more, testing their data efficiency and scalability for long-horizon MoMa tasks.
The ablation results, summarized in Figure 2, show a positive upward trend in performance for RT-1 (represented by the pink line) as the training data increases, albeit a weak one. This trend is evident in its average rate of change (RoC). This shows that scaling the data does improve RT-1âs overall performance. A similar trend is observed in the language, vision, and multimodal communities [39, 40, 41], making it encouraging to see this pattern emerge in the context of long-horizon MoMa tasks as well. Our results reflect that these BC models are data inefficient for long-horizon MoMa tasks, as the fewer the data, the worse it generally performs. This raises questions about the source of inefficiency, such as model architecture, explored in the following section.
Impact of Architecture. To assess whether data inefficiency stems from the architecture, we evaluate and compare different architectures. While MG-MoMa was already benchmarked and provides some architectural comparison with RT-1, it is not sufficiently different since it also uses a transformer [31]. To explore distinct alternatives, we replace RT-1âs transformer with alternate architectures while keeping all other components, such as encoders, unchanged to ensure a controlled comparison. Specifically, we replace the transformer with an LSTM [42], which we refer to as RL-1. Likewise, we incorporate the Mamba architecture [43], leveraging its ability to handle long sequences [43], and call this RM-1. To understand how these alternative architectures perform with increasing data, we evaluate both RL-1 and RM-1 under the Task Generalization setup using the same subset-based evaluation as previously conducted. By comparing their trends/RoCs and full dataset results against RT-1âs, we aim to determine whether data inefficiency is tied to architectural limitations or not.
As shown in Figure 2, both RL-1 and RM-1 exhibit upward trends, indicating that they did not learn more efficiently than RT-1 and that the data inefficiency is spanning across different architectures. Identifying the exact causes of data inefficiency requires further investigation, which is beyond the scope of this work and is left for the community to explore in the future.